please help me to solve this..
find the maximum and minimum values of f(x)=x^3-6x^2+24x+4

Answer by Alan3354(69443)

please help me to solve this..
find the maximum and minimum values of f(x)=x^3-6x^2+24x+4
----------------
f(x)=x^3-6x^2+24x+4
f'(x)= 3x^2 - 12x + 24 = 0
--> x = 2 ± 2i
No max or min
-------------